OpenStudy (anonymous):

Solve the equation. Show work plz 0.25 r – 0.125 + 0.5 r = 0.5 + r

OpenStudy (anonymous):

0.75r - 0.125 = 0.5 + r -0.5 - 0.125 = r - 0.75 - 0.625 = 0.25r Therefore, r = -2.5

OpenStudy (anonymous):

in the second step it is -0.75r. =P

OpenStudy (anonymous):

0.25 r - 0.125 + 0.5 r = 0.5 + r\[\frac{25}{100}r-\frac{125}{1000}+\frac{5}{10}r=\frac{5}{10}+r \]\[\frac{r}{4}-\frac{1}{8}+\frac{r}{2}=\frac{1}{2}+r \]\[-5-2 r=0\]\[r=-\frac{5}{2} \]